# Encoder vs Decoder: Comparative Analysis of Encoder and Decoder Language Models on Multilingual NLU Tasks

### v1

@@ Abstract
This paper explores the performance of encoder and decoder language models on multilingual Natural Language Understanding (NLU) tasks, with a broad focus on Germanic languages. Building upon the ScandEval benchmark, which initially was restricted to evaluating encoder models, we extend the evaluation framework to include decoder models. We introduce a method for evaluating decoder models on NLU tasks and apply it to the languages Danish, Swedish, Norwegian, Icelandic, Faroese, German, Dutch, and English. Through a series of experiments and analyses, we address key research questions regarding the comparative performance of encoder and decoder models, the impact of NLU task types, and the variation across language resources. Our findings reveal that decoder models can achieve significantly better NLU performance than encoder models, with nuances observed across different tasks and languages. Additionally, we investigate the correlation between decoders and task performance via a UMAP analysis, shedding light on the unique capabilities of decoder and encoder models. This study contributes to a deeper understanding of language model paradigms in NLU tasks and provides valuable insights for model selection and evaluation in multilingual settings.
Encoder vs Decoder: Comparative Analysis of Encoder and Decoder Language Models on Multilingual NLU Tasks
Dan Saattrup Nielsen The Alexandra Institute dan.nielsen@alexandra.dk                        Kenneth Enevoldsen University of Aarhus kenneth.enevoldsen@cas.au.dk                        Peter Schneider-Kamp University of Southern Denmark petersk@imada.sdu.dk
@@ 1 Introduction
    Can state-of-the-art decoder models achieve significantly better NLU performance than state-of-the-art finetuned encoder models?
@@
    We extend the languages supported by the ScandEval benchmarking framework by German, Dutch and English. Together with Danish, Swedish, Norwegian, Icelandic and Faroese, ScandEval now provides virtually full coverage of the Germanic language group.
@@
    We give a positive answer to (Q1), showing that decoder models can achieve significantly better NLU performance than encoder models. This depends on the language in question however, giving a partially positive answer to (Q3).
@@ 2.1 Comparing Encoder and Decoder Models
There has been a number of studies in recent years comparing encoder models to decoder models. Zhong et al. (2023) compared GPT-3.5-turbo (January 2023 version) to (finetuned versions of) the base- and large versions of BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019) on the English GLUE benchmark Wang et al. (2018). They found that GPT-3.5-turbo was on average on par with the base-sized encoder models, but fell short of the large-sized ones. They also note that despite being on par with the base-sized models, there was a big discrepancy between the models on individual tasks, with GPT-3.5-turbo for instance being better on the inference tasks while being worse on the paraphrase tasks. We note however that they only evaluate the decoder model in a zero-shot setting, and furthermore they only evaluate the models on 25 samples for each class in the development split, leading to a potential lack of robustness in their evaluation.
Wang et al. (2023) compared GPT-3.5-turbo (January 2023 version) to a finetuned version of the base-sized BERT model on 18 English benchmark datasets related to sentiment analysis. Like Zhong et al. (2023), they found that the zero-shot performance of GPT-3.5-turbo was on par with the base-sized BERT model, and that the few-shot performance of GPT-3.5-turbo (with 27 few-shot examples) was slightly better than BERT, on average. Their test sets contained, on average, 538 samples, which is a significant improvement over Zhong et al. (2023). However, the narrow focus on the evaluation tasks as well as only benchmarking a single encoder and decoder model makes it hard to generalise the results to other tasks and models.
Kocoń et al. (2023) built a benchmark suite of 25 tasks, where 21 of these tasks were classification tasks (binary, multi-class and multi-label), 3 being question answering tasks and the last one being a token classification task. Two of the classification tasks were in Polish, the rest being English. They compared the zero-shot and few-shot performance of GPT-3.5-turbo (January 2023 version) to the state-of-the-art encoder performance on each task. GPT-3.5-turbo was generally found to be worse than state-of-the-art encoder models. They also evaluated GPT-4 on five of the tasks (inference, question-answering and emotion datasets), and only found GPT-4 to be marginally better than GPT-3.5-turbo, still far off the encoder models.
Qiu and Jin (2024) compared GPT-3.5-turbo (January 2023 version) to a finetuned version of the base-sized BERT model on three manually curated English multi-class classification datasets with 19, 12 and 7 test samples, respectively, where they found that the BERT model performed marginally better than GPT-3.5-turbo in a few-shot setting (and that the zero-shot performance was significantly worse). The tiny test sets make it hard to generalise the results, however.
@@ 2.2 Benchmarks of Generative Language Models
There has been several language-specific benchmarks introduced as well. NorBench Samuel et al. (2023) is a collection of Norwegian evaluation datasets moreso than a dedicated evaluation framework. Further, several datasets in this collection (NorQuAD, NoReC and NorNE) are already part of ScandEval. SuperLim Berdičevskis et al. (2023) falls into the same category for Swedish. DUMB de Vries et al. (2023) is a Dutch benchmarking framework, which is only focused on encoder models. Danoliterate Holm (2024) is a Danish benchmarking framework, which has been based largely on the present ScandEval benchmark, and solely focused on evaluating decoder models. Aside from language modelling performance, the Danoliterate benchmark also measures calibration, efficiency, toxicity and fairness. While the development of language-specific benchmarks is important, it leads to too little overview of trends across benchmarks and languages and incentivises model development focused on monolingual models ignoring a potential broader appeal. ScandEval provides a unified and robust approach for comparison across model categories and Germanic languages.
@@ 3 Datasets
3.1 Named Entity Recognition
3.2 Sentiment Classification
3.3 Linguistic Acceptability
3.4 Extractive Question Answering
Here we use the ScandiQA dataset Nielsen (2023) for Danish and Swedish, but replace the manually translated Norwegian ScandiQA dataset with the new curated NorQuAD dataset Ivanova et al. (2023). We further add the new Natural Questions in Icelandic dataset Snæbjarnarson and Einarsson (2022) for Icelandic. For German and English we add the existing extractive question-answering datasets GermanQuAD Möller et al. (2021) and SQuAD Rajpurkar et al. (2016), respectively. For Dutch we add the machine translated version of SQuAD to Dutch Havinga (2023). We acknowledge that a machine translated dataset is not ideal, and will return to this point when analysing the evaluation results.
@@ 3.4 Extractive Question Answering
      Angry Tweets Pauli et al. (2021) Danish 1,024 256 2,048 3
      SweReC Svensson (2017) Swedish 1,024 256 2,048 3
      NoReC Velldal et al. (2018) Norwegian 1,024 256 2,048 3
      SB10k Cieliebak et al. (2017) German 1,024 256 1,024 3
      Dutch Social Gupta (2022) Dutch 1,024 256 1,024 3
      SST5 Socher et al. (2013) English 1,024 256 2,048 3
@@
      ScaLA-da Nielsen (2023) Danish 1,024 256 2,048 3
      ScaLA-sv Nielsen (2023) Swedish 1,024 256 2,048 3
      ScaLA-nb Nielsen (2023) Norwegian Bokmål 1,024 256 2,048 3
      ScaLA-nn Nielsen (2023) Norwegian Nynorsk 1,024 256 2,048 3
      ScaLA-is Nielsen (2023) Icelandic 1,024 256 2,048 3
      ScaLA-fo Nielsen (2023) Faroese 1,024 256 1,024 3
      ScaLA-de Nielsen (2023) German 1,024 256 2,048 3
      ScaLA-nl Nielsen (2023) Dutch 1,024 256 2,048 3
      ScaLA-en Nielsen (2023) English 1,024 256 2,048 3
@@ 4.1 Formulating NLU Tasks as Generative Tasks
[label prefix]: [label]
[document prefix]: [document]
@@
We found that the separation of the few-shot examples with double newlines makes it easier to know when to stop the generation - for the same reason, we ensure that there are no double newlines in any of the documents. See the prompts used for the English datasets in Table 4.1; a full table of the prompts used for all the tasks in all the languages can be found in the appendix.
@@
Table 2: The English prompt templates used for the datasets. See all the prompt templates in the appendix.
@@ 4.3 Score Aggregation Method
We then add this to the rank score of the first model. We continue this process for all the models to get the rank scores for the dataset, and to compute the overall score for the model, we take the mean of the rank scores for the datasets. An overview of this aggregation method can be found in Algorithm 1.
We note that the mean rank score has an intuitive interpretation: it is the average number of standard deviations from the best scoring model (+1).
@@
Algorithm 1 The mean rank score algorithm.
2:Output: The mean rank score of each model
3:$R\leftarrow\textrm{matrix of size }M\times D$
4:for $\hat{d}\in D$ do
6:     for $\hat{m}\in\textrm{sortedModels}$ do
7:         if $\hat{m}$ is the first model then
8:              $\rho\leftarrow 1$
9:              $\textrm{previousScores}\leftarrow S_{\hat{m},\hat{d}}$
10:         else
[…]
12:              if $\textrm{pValue}<0.05$ then
15:                  $\textrm{previousScores}\leftarrow S_{\hat{m},\hat{d}}$
16:              end if
17:         end if
18:         $R_{\hat{m},\hat{d}}\leftarrow\rho$
19:     end for
20:end for
21:$\textrm{output}\leftarrow\nicefrac{{1}}{{D}}\cdot\Sigma_{d}R_{m,d}$
@@ 5.1 Comparative Performance Analysis on High- and Low-resource Languages
Excerpts of the English, Danish and Icelandic leaderboards can be found in Table 3, Table 4 and Table 5, respectively. We found that these three represent three main categories of languages with respect to the open-closed source divide. Similar excerpts for the remaining languages (Swedish, Norwegian, Faroese, German and Dutch) can be found in the appendix. The full leaderboards for all the languages can be found at https://scandeval.com.
From the English results we see that the state-of-the-art decoder model GPT-4-0613 Achiam et al. (2023) is still outperformed by the DeBERTa-v3-large and DeBERTa-v3-base models He et al. (2020) as well as the ELECTRA-base model Clark et al. (2020). Here GPT-4-0613 is, on average, 0.44 standard deviations worse than the best model. The same pattern is seen for Norwegian, Dutch, German and Faroese; see the appendix for the corresponding leaderboard excerpts.
In contrast, on the Danish leaderboard, the top-3 models are all decoder models, with GPT-4-0613 and GPT-4-1106-preview OpenAI (2023b) in the lead, followed by the closed-source DanskGPT-Chat-Llama3-70B model from Syv.AI [^1] https://www.syv.ai/, being a continuation of the Llama-3-70B model AI@Meta (2024). The GPT-4-0613 model is, on average, 0.24 standard deviations from the best model. Similar results were found with Swedish; see the appendix for the corresponding leaderboard excerpt.
@@
We can thus give an affirmative answer to research question (Q1), showing that decoder models can achieve significantly better NLU performance than encoder models, albeit with an order of magnitude more model parameters. For (Q3), we see that this varies between languages, but without being correlated to the language resource spectrum.
@@ 5.2 Task Analysis
Firstly, we analyse whether the score distribution across the four NLU tasks is different for the encoder and decoder models. This is done by applying a UMAP McInnes et al. (2018) to the results of a given leaderboard, stratified by both the mean rank score and whether the model is generative. UMAP plots for the English, Danish, Swedish, Norwegian, German and Dutch leaderboards can be found in Figure 1.
We see that the worst and best performing models have similar distributions, irrespective of whether they are generative or not. However, we also note that the rest of encoder and decoder models follow different “paths” in the UMAP space, leading to our hypothesis that they learn different things during their development.
@@
In Figure 2 we show the correlation between a model being generative and its performance on the four NLU tasks. We see that being generative is a strong predictor for good question answering performance, as well as poor named entity recognition and linguistic acceptability performance. The correlation is weaker for sentiment classification and varies across languages. We also see that these findings seem to generalise across languages, both high- and low-resource. The large question answering performance persists for non-instruction-tuned decoder models (see the leaderboards at https://scandeval.com), showing a likely side-effect of the pre-training algorithm or the architecture of decoder models making them better at this task.
@@ 7 Conclusion
We have extended the ScandEval benchmark to include the evaluation of decoder models, as well as including three new languages: German, Dutch and English. From the analysis of the corresponding results we found that decoder models can achieve significantly better NLU performance than encoder models, but that this varies between languages. We have also shown that being generative is strongly correlated with both good question answering performance and poor performance for named entity recognition and linguistic acceptability. Our analysis showed that the “path” from the worst to the best-performing models in the UMAP space is different for encoder and decoder models, indicating that they learn different things during their development.
@@
We have taken care to ensure that the evaluation is fair and unbiased, both through our selection of the datasets in the benchmark as well as through our choice of aggregation method of the scores.

### v2

@@ Abstract
This paper explores the performance of encoder and decoder language models on multilingual Natural Language Understanding (NLU) tasks, with a broad focus on Germanic languages. Building upon the ScandEval benchmark, initially restricted to evaluating encoder models, we extend the evaluation framework to include decoder models. We introduce a method for evaluating decoder models on NLU tasks and apply it to the languages Danish, Swedish, Norwegian, Icelandic, Faroese, German, Dutch, and English. Through a series of experiments and analyses, we also address research questions regarding the comparative performance of encoder and decoder models, the impact of NLU task types, and the variation across language resources. Our findings reveal that encoder models can achieve significantly better NLU performance than decoder models despite having orders of magnitude fewer parameters. Additionally, we investigate the correlation between decoders and task performance via a UMAP analysis, shedding light on the unique capabilities of decoder and encoder models. This study contributes to a deeper understanding of language model paradigms in NLU tasks and provides valuable insights for model selection and evaluation in multilingual settings.
@@ 1 Introduction
    Can state-of-the-art finetuned encoder models achieve significantly better NLU performance than state-of-the-art decoder models?
@@
    We extend the languages supported by the ScandEval benchmarking framework by German, Dutch and English. Together with Danish, Swedish, Norwegian, Icelandic and Faroese, ScandEval now provides coverage of all Germanic languages except Afrikaans and the Frisian languages.
@@
    We give a positive answer to (Q1), showing that encoder models achieve significantly better NLU performance than encoder models in several languages. This depends on the language in question however, giving a partially positive answer to (Q3).
@@ 2.1 Comparing Encoder and Decoder Models
There has been a number of studies in recent years comparing encoder models to decoder models. Zhong et al. (2023) compared GPT-3.5-turbo (January 2023 version) to (finetuned versions of) the base and large versions of BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019) on the English GLUE benchmark Wang et al. (2018). They find that GPT-3.5-turbo is on average on par with the base-sized encoder models, but falls short of the large-sized ones. They also note that despite being on par with the base-sized models, there is a big discrepancy between the models on individual tasks, with GPT-3.5-turbo for instance being better on the inference tasks while being worse on the paraphrase tasks. We note however that they only evaluate the decoder model in a zero-shot setting, and furthermore they only evaluate the models on 25 samples for each class in the development split, leading to a potential lack of robustness in their evaluation.
Wang et al. (2023) compares GPT-3.5-turbo (January 2023 version) to a finetuned version of the base-sized BERT model on 18 English benchmark datasets related to sentiment analysis. Like Zhong et al. (2023), they find that the zero-shot performance of GPT-3.5-turbo is on par with the base-sized BERT model, and that the few-shot performance of GPT-3.5-turbo (with 27 few-shot examples) is slightly better than BERT, on average. Their test sets contained, on average, 538 samples, which is a significant improvement over Zhong et al. (2023). However, the narrow focus on the evaluation tasks as well as only benchmarking a single encoder and decoder model makes it hard to generalise the results to other tasks and models.
Kocoń et al. (2023) built a benchmark suite of 25 tasks, where 21 of these tasks are classification tasks (binary, multi-class and multi-label), 3 being question answering tasks and the last one being a token classification task. Two of the classification tasks are in Polish and the rest in English. They compare the zero-shot and few-shot performance of GPT-3.5-turbo (January 2023 version) to the state-of-the-art encoder performance on each task. GPT-3.5-turbo is generally found to be worse than state-of-the-art encoder models. They also evaluate GPT-4 on five of the tasks (inference, question-answering and emotion datasets), and only find GPT-4 to be marginally better than GPT-3.5-turbo, still far off the encoder models.
Qiu and Jin (2024) compare GPT-3.5-turbo (January 2023 version) to a finetuned version of the base-sized BERT model on three manually curated English multi-class classification datasets with 19, 12 and 7 test samples, respectively, where they find that the BERT model performs marginally better than GPT-3.5-turbo in a few-shot setting (and that the zero-shot performance is significantly worse). The tiny test sets make it hard to generalise the results, however.
@@ 2.2 Benchmarks of Generative Language Models
There has been several language-specific benchmarks introduced as well. NorBench Samuel et al. (2023) is a collection of Norwegian evaluation datasets moreso than a dedicated evaluation framework. Further, several datasets in this collection (NorQuAD, NoReC and NorNE) are already part of ScandEval. SuperLim Berdičevskis et al. (2023) falls into the same category for Swedish. DUMB de Vries et al. (2023) is a Dutch benchmarking framework, which is only focused on encoder models. Danoliterate Holm (2024) is a Danish benchmarking framework which is solely focused on evaluating decoder models, and whose datasets largely overlap with the Danish datasets in ScandEval, albeit with a different evaluation methodology. Aside from language modelling performance, the Danoliterate benchmark also measures calibration, efficiency, toxicity and fairness. While the development of language-specific benchmarks is important, it leads to too little overview of trends across benchmarks and languages and incentivises model development focused on monolingual models ignoring a potential broader appeal. ScandEval provides a unified and robust approach for comparison across model categories and Germanic languages.
@@ 3 Datasets
      Angry Tweets Pauli et al. (2021) Danish 1,024 256 2,048 12
      SweReC Svensson (2017) Swedish 1,024 256 2,048 12
      NoReC Velldal et al. (2018) Norwegian 1,024 256 2,048 12
      SB10k Cieliebak et al. (2017) German 1,024 256 1,024 12
      Dutch Social Gupta (2022) Dutch 1,024 256 1,024 12
      SST5 Socher et al. (2013) English 1,024 256 2,048 12
@@
      ScaLA-da Nielsen (2023) Danish 1,024 256 2,048 12
      ScaLA-sv Nielsen (2023) Swedish 1,024 256 2,048 12
      ScaLA-nb Nielsen (2023) Norwegian Bokmål 1,024 256 2,048 12
      ScaLA-nn Nielsen (2023) Norwegian Nynorsk 1,024 256 2,048 12
      ScaLA-is Nielsen (2023) Icelandic 1,024 256 2,048 12
      ScaLA-fo Nielsen (2023) Faroese 1,024 256 1,024 12
      ScaLA-de Nielsen (2023) German 1,024 256 2,048 12
      ScaLA-nl Nielsen (2023) Dutch 1,024 256 2,048 12
      ScaLA-en Nielsen (2023) English 1,024 256 2,048 12
@@
3.1 Named Entity Recognition
3.2 Sentiment Classification
3.3 Linguistic Acceptability
3.4 Extractive Question Answering
Here we use the ScandiQA dataset Nielsen (2023) for Danish and Swedish, but replace the manually translated Norwegian ScandiQA dataset with the new curated NorQuAD dataset Ivanova et al. (2023). We further add the new Natural Questions in Icelandic dataset Snæbjarnarson and Einarsson (2022) for Icelandic. For German and English we add the existing extractive question-answering datasets GermanQuAD Möller et al. (2021) and SQuAD Rajpurkar et al. (2016), respectively. For Dutch we add the machine translated version of SQuAD to Dutch Havinga (2023).
@@ 4.1 Formulating NLU Tasks as Generative Tasks
We found that the separation of the few-shot examples with double newlines makes it easier to know when to stop the generation - for the same reason, we ensure that there are no double newlines in any of the documents. See the prompts used for the English datasets in Table 4.1; a full table of the prompts used for all the tasks in all the languages can be found in Appendix LABEL:apx:datasets.
@@
Table 2: The English prompt templates used for the datasets. See all the prompt templates in Appendix LABEL:apx:datasets.
@@ 4.2 Evaluation Methodology
The number of few-shot examples for each dataset was determined on a heuristic basis, where we wanted to include as many examples as possible, while making sure that the token count was sufficiently low to not bias the evaluation towards models with a longer context length. All the NER, sentiment classification and linguistic acceptability datasets have prompt sizes around 1,000 tokens with the Mistral-7B-v0.1 tokeniser Jiang et al. (2023), with the question answering datasets having around 2,000 tokens. This is also the reason for the discrepancy with the NorQuAD dataset, as the samples are much longer than the other question answering datasets.
@@ 4.3 Score Aggregation Method
We then add this to the rank score of the first model. We continue this process for all the models to get the rank scores for the dataset, and to compute the overall score for the model, we take the mean of the rank scores for the datasets. An overview of this aggregation method can be found in Appendix B. We note that the mean rank score has an intuitive interpretation: it is the average number of standard deviations from the best scoring model (+1).
@@ 5.1 Comparative Performance Analysis on High- and Low-resource Languages
Excerpts of the English, Danish and Icelandic leaderboards can be found in Table 3, Table 4 and Table 5, respectively. We found that these three represent three main categories of languages with respect to the open-closed source divide. Similar excerpts for the remaining languages (Swedish, Norwegian, Faroese, German and Dutch) can be found in Appendix C. The full leaderboards for all the languages can be found at https://scandeval.com.
From the English results we see that the state-of-the-art decoder model GPT-4-0613 Achiam et al. (2023) is still outperformed by the DeBERTa-v3-large and DeBERTa-v3-base models He et al. (2020) as well as the ELECTRA-base model Clark et al. (2020). Here GPT-4-0613 is, on average, 0.44 standard deviations worse than the best model. The same pattern is seen for Norwegian, Dutch, German and Faroese; see Appendix C for the corresponding leaderboard excerpts.
In contrast, on the Danish leaderboard, the top-3 models are all decoder models, with GPT-4-0613 and GPT-4-1106-preview OpenAI (2023b) in the lead, followed by the closed-source DanskGPT-Chat-Llama3-70B model from Syv.AI [^1] https://www.syv.ai/, being a continuation of the Llama-3-70B model AI@Meta (2024). The GPT-4-0613 model is, on average, 0.24 standard deviations from the best model. Similar results were found with Swedish; see Appendix C for the corresponding leaderboard excerpt.
@@
We can thus give an affirmative answer to research question (Q1), showing that encoder models can achieve significantly better NLU performance than decoder models, even though they have an order of magnitude fewer model parameters. For (Q3), we see that this varies between languages, but without being correlated to the language resource spectrum.
@@ 5.2 Task Analysis
Firstly, we analyse whether the score distribution across the four NLU tasks is different for the encoder and decoder models. This is done by applying a UMAP McInnes et al. (2018) to the results of a given leaderboard, which is a dimensionality reduction method that both takes into account the global and local structure of the underlying data - it can thus be viewed as a middle ground between a principal component analysis Pearson (1901) and a t-distributed stochastic neighbour embedding Hinton and Roweis (2002). The resulting reduction thus contains a single two-dimensional representation of each model. UMAP plots for the English, Danish, Swedish, Norwegian, German and Dutch leaderboards can be found in Figure 1, where we also mark the mean rank score for each model, as well as whether the model is generative.
We see that the worst and best performing models have similar distributions, irrespective of whether they are generative or not. However, we also note that the rest of encoder and decoder models follow different “paths” in the UMAP space, leading to our hypothesis that the different architectures have different task preferences.
@@
In Figure 2 we show the correlation between a model being generative and its performance on the four NLU tasks. We see that being generative is a strong predictor for good question answering performance, as well as poor named entity recognition and linguistic acceptability performance. The correlation is weaker for sentiment classification and varies across languages. We also see that these findings seem to generalise across languages, both high- and low-resource. The large question answering performance persists for non-instruction-tuned decoder models (see the leaderboards at https://scandeval.com), showing a likely side-effect of the pre-training algorithm or the architecture of decoder models making them better at this task. We also note that generative models perform substantially better at the English sentiment classification dataset SST5 compared to the other sentiment classification datasets - we will return to this in the discussion.
@@ 6 Discussion
Some of the datasets in the benchmark are translations of American datasets, which we acknowledge is not ideal and encourage the development of gold-standard replacements of these. This concerns the Dutch question answering dataset, which is machine translated, as well as the Danish and Swedish question answering datasets, where the questions and answers have been manually translated. Manual translations are typically better than machine translations, but it nevertheless means that the content is biased towards questions pertinent to the American context. Some datasets are furthermore missing. This concerns Icelandic and Faroese sentiment analysis, as well as Faroese question answering. Efforts are currently underway to remedy this.
Lastly, we note that the English sentiment classification dataset SST5 is the only dataset where generative models perform substantially better than encoder models. We speculate that this is either due to the dataset simply being significantly easier than the others, or that the test data has leaked into the pretraining datasets of the generative models. The dataset is part of the FLAN collection Wei et al. , which is for instance included in the Dolma dataset Soldaini et al. (2024), which is used to pretrain the OLMo model Groeneveld et al. (2024), being one of the generative models that is performing very well on this dataset. Leakage is therefore possible, and we encourage the community to investigate this further.
@@ 7 Conclusion
We have extended the ScandEval benchmark to include the evaluation of decoder models, as well as including three new languages: German, Dutch and English. From the analysis of the corresponding results we found that encoder models can achieve significantly better NLU performance than decoder models despite having orders of magnitude fewer parameters, but that this varies between languages. We have also shown that being generative is strongly correlated with both good question answering performance and poor performance for named entity recognition and linguistic acceptability. Our analysis showed that the “path” from the worst to the best-performing models in the UMAP space is different for encoder and decoder models, indicating an architecture-specific task-preference.
@@
We have made efforts towards making the evaluation as fair and unbiased as possible, both through our selection of the datasets in the benchmark as well as through our choice of aggregation method of the scores. However, we have not conducted extensive bias analyses on the individual evaluation datasets.